3) Find an equation, y = mx + b, of the line shown below. 3 points 24. OD TBS
Answer #1

We have... given 03,0) 5 -4 -3 2 1 2 3 4 5 6 7 21 -2). to we have y = mx + b find an equation from above line Here m= slope om = 92-9 X2 - - 2 - 0 0 - 3 - 2 2 from above gruph we see that the line crosses the y-axis at -2 so the y intercept is -2 equ